OpenStudy (anonymous):

Which of the following is the conjugate of a complex number with 2 as the real part and -8 as the imaginary part? a. -2 + 8i b. 2 + 8i c. 2 - 8i d. -2 - 8i

OpenStudy (anonymous):

@GamerGirl.1

OpenStudy (anonymous):

complex number=2-8i its conjugate =2+8i